OpenStudy (anonymous):

: Set-up a system AND solve. (60 points) Suppose you decide you want to live a healthier lifestyle. A new fitness club has opened. At the time when the new club opened, there were 100 people that pre-registered and each week an average of 7 new members joined. When the local YMCA opened, there were 573 members and enrollment has decreased by 3 members since the new fitness club opened. After how many weeks will the total number of members for each fitness club be the same? Equation 1: Equation 2 After how many weeks will the total number of members for each fitness club be the same?

OpenStudy (anonymous):

the equations are Equation 1: b+ s = 60 Equation 2: 10b + 12s = 500

OpenStudy (anonymous):

100 + 7x = 573 - 3x ?

OpenStudy (anonymous):

music101 is right... solve for x with that equation, then x will be the number of weeks that has passed to when each have the same number of members. I do not know where you came up with those equations you have though.